The second angle of a triangle measures twice as large as the first. If the third angle measures 40 º more than the first, find the measures of all three angles.
Answer by checkley79(3341)

You have x,2x (x+40) for the 3 angles.
x+2x+x+40=180
4x=180-40
4x=140
x=140/4
x=35 degrees for tghe first angle
2*35=70 degrees for the second angle.
35+40=75 degrees for the third angle.
Proof:
35+70+75=180
180=180